# Combinatorial Optimization using Comparison Oracles

### v1

@@ C.3.2 Stronger Oracles
C.4 Illustrative Examples
C.4.1 Separations between Comparison and Value Queries
@@
C.4.2 Non-Reconstructable Graphs
The unweighted graph pairs $(K_{2},\bar{K_{2}})$ and $(K_{3},\bar{K_{3}})$ are not distinguishable, since essentially the only feasible sets $S$ are singleton sets (recall that by symmetry, $f(S)=f(\bar{S})$), and all of them have the same cut value. However, as we showed in Theorem˜1.6, we can recover all unweighted simple graphs other than $K_{2},\bar{K_{2}},K_{3},\bar{K_{3}}$ using comparison queries.
@@ C.4.2 Non-Reconstructable Graphs
C.4.3 Finding Heaviest Edge Incident to a Vertex
@@
Appendix D Matroids, Matchings, and Paths
@@ D.1 Matroid Bases
Theorem D.1 (Matroid Bases).
@@
For the first part, let $B$ be the basis that extends the independent set $C-e^{\prime}$. This means that the fundamental circuit in $B+e^{\prime}$ is $C$ implying that $B^{\prime}:=B+e^{\prime}-e$ is also a basis. The second part is proved in [Oxl11, Proposition 4.1.2]. The equivalence classes in lemma˜D.2(2) are called the connected components of $\mathcal{M}$. ∎
@@
Given Lemma˜D.3, the matroid decompostion step requires $O(n^{2})$ matroid independence calls (in fact, the calls are always about whether a set is a basis or not) to build the exchange graph using Lemma˜D.3 part 1 and $O(n^{2})$ time to identify connected components.
Next, to simulate any pairwise comparison between elements $x,y$, we obtain the two bases $B\ni x,B^{\prime}\ni y$ such that $B\Delta B^{\prime}=\{x,y\}$ using part 2 of Lemma˜D.3. The implementation of Kruskal’s algorithm requires sorting the elements in each connected component which requires $O(c_{i}\log c_{i})$ comparisons in each component with $c_{i}$ elements. Each such comparison takes $O(n)$ time to find the path and write down the bases to input to the comparison oracle. Adding everything, we get the $O(n^{2}\log n)$ time bound as claimed. ∎
D.2 Matroid Intersection
@@
Theorem D.4 (Matroid Intersection).
@@
By iteratively applying Lemma˜D.5, we construct extremal matchings of increasing size, from size 0 to $n/2$, and return the one with minimum total weight. Moreover, to implement this using only comparisons between matchings, we describe a variant of the Bellman-Ford algorithm. For each $u\in L$, let $p_{u}^{(k)}$ be the shortest $M$-alternating path from $L^{\prime}$ to $u$ using at most $k$ matching edges. For matched $u\in L$, let $u^{\prime}\in R$ be its partner. We initialize:
@@
D.3 Generalization to Matroid Intersection
@@
Lemma D.6 (Matching lemmas).
@@
Lemma D.9 (Minimal paths and cycles).
@@
Consider the multiset union of edges $N\cup N^{\prime}\cup 2M$ which takes two copies of every edge from $M$ and a copy of every edge from $N$ and $N^{\prime}$. Since these set of edges enter and leave every vertex exactly twice, the set of edges can be decomposed into a collection of cycles $C_{1},\dots,C_{p}$ for some $p\geq 2$ such that some of the cycles have strictly fewer vertices than $C$. We also have $\sum_{i\in p}\ell(C_{i})=2\ell(C)$. Using Lemma˜D.8(i), we know $\ell(C_{i})\geq 0$ and since $p\geq 2$, there exists a cycle $C_{i}$ with lesser weight or the same weight with lesser number of vertices as $C$ contradicting the minimality of $C$.
@@
    (No re-entry) First, we prove that a minimal path in $\mathcal{P}(y,t)$ never re-enters $X_{1}$ again. If $P=(x_{0},y_{1},x_{1},\dots,y_{k},x_{k},\dots,x_{t-1},y_{t})$ with $x_{k}\in X_{1}$, then observe that there is an edge from $y_{k}$ to $x_{0}$ because $Y+x_{0}\in\mathcal{I}_{1}$ as $x_{0}\in\mathcal{I}_{1}$ which implies $Y+x_{0}-y_{k}\in\mathcal{I}_{1}$ and hence the edge. If we let $C=(x_{0},y_{1},x_{1},\dots,y_{k},x_{0})$, and $P^{\prime}=(x_{k},y_{k+1},\dots,x_{t-1},y_{t})$, we have $\ell(P)=\ell(C)+\ell(P^{\prime})$ implying that $P^{\prime}$ has smaller length and lesser number of arcs than $P$ as $\ell(C)\geq 0$ from Lemma˜D.8(i) contradicting the minimality of $P$. An identical argument can be used to show that vertices are not revisited again in minimal paths.
@@
    Since these set of edges enter and leave every vertex exactly twice, the set of edges can be decomposed into a collection of cycles $C_{1},\dots,C_{p}$ for some $p\geq 2$ such that some of the cycles have strictly fewer vertices than $C$. We also have $\sum\limits_{i\in p}\ell(C_{i})=2\ell(C)$. Using Lemma˜D.8(i), we know $\ell(C_{i})\geq 0$. Consider the two different cycles (say) $C_{1},C_{2}$ that each contain a copy of the edge $(y_{t},x_{0})$. We know that one of these cycles (say) $C_{1}$ should have $\ell(C_{1})\leq\ell(C)$ and $|C_{1}|<|C|$. Removing the $(y_{t},x_{0})$ edge from $C_{1}$ gives a path (say) $P_{1}$ such that $(\ell(P),|P|)\succeq(\ell(P_{1}),|P_{1}|)$ contradicting the minimality of $P$.
@@
Lemma˜D.9 allows us to compare lengths of minimal paths using only comparisons of matroid intersections:
@@
Proof of Theorem˜D.4.
Let $Y_{t}$ be the extreme common independent set of size exactly $t$ starting with $Y_{0}=\emptyset$. Use the modified Bellman-Ford Algorithm to find the shortest augmenting path $P_{t}$ in the exchange graph $H(\mathcal{M}_{1},\mathcal{M}_{2},Y_{t}))$. Using Lemma˜D.8 part (i), the exchange graph has no negative cycles. The shortest path computation takes $O(n^{3})$ comparison queries and time (see section˜D.4). After every update, we know that $Y_{t+1}:=Y_{t}\Delta P_{t}$ is an extreme set using Lemma˜D.8 part (ii). We continue for $n$ steps as long as there is a directed path from $X_{1}$ to $X_{2}$ in the exchange graph. Using Lemma˜D.7, we know that the maximum cardinality intersection is reached when an augmenting path is not present anymore. Finally, we can compare the weights of all the extreme sets obtained to output the minimum weight common intersection $O(n^{4})$ comparisons and $O(n^{4})$ time. ∎
D.4 Shortest paths

### v2

@@ D.3.2 Stronger Oracles
D.4 Illustrative Examples
D.4.1 Separations between Comparison and Value Queries
@@
D.4.2 Non-Reconstructable Graphs
The unweighted graph pairs $(K_{2},\bar{K_{2}})$ and $(K_{3},\bar{K_{3}})$ are not distinguishable, since essentially the only feasible sets $S$ are singleton sets (recall that by symmetry, $f(S)=f(\bar{S})$), and all of them have the same cut value. However, as we showed in Theorem˜1.7, we can recover all unweighted simple graphs other than $K_{2},\bar{K_{2}},K_{3},\bar{K_{3}}$ using comparison queries.
@@ D.4.2 Non-Reconstructable Graphs
D.4.3 Finding Heaviest Edge Incident to a Vertex
@@
Appendix E Matroids, Matchings, and Paths
@@ E.1 Matroid Bases
Theorem E.1 (Matroid Bases).
@@
For the first part, let $B$ be the basis that extends the independent set $C-e^{\prime}$. This means that the fundamental circuit in $B+e^{\prime}$ is $C$ implying that $B^{\prime}:=B+e^{\prime}-e$ is also a basis. The second part is proved in [Oxl11, Proposition 4.1.2]. The equivalence classes in lemma˜E.2(2) are called the connected components of $\mathcal{M}$. ∎
@@
Given Lemma˜E.3, the matroid decompostion step requires $O(n^{2})$ matroid independence calls (in fact, the calls are always about whether a set is a basis or not) to build the exchange graph using Lemma˜E.3 part 1 and $O(n^{2})$ time to identify connected components.
Next, to simulate any pairwise comparison between elements $x,y$, we obtain the two bases $B\ni x,B^{\prime}\ni y$ such that $B\Delta B^{\prime}=\{x,y\}$ using part 2 of Lemma˜E.3. The implementation of Kruskal’s algorithm requires sorting the elements in each connected component which requires $O(c_{i}\log c_{i})$ comparisons in each component with $c_{i}$ elements. Each such comparison takes $O(n)$ time to find the path and write down the bases to input to the comparison oracle. Adding everything, we get the $O(n^{2}\log n)$ time bound as claimed. ∎
E.2 Matroid Intersection
@@
Theorem E.4 (Matroid Intersection).
@@
By iteratively applying Lemma˜E.5, we construct extremal matchings of increasing size, from size 0 to $n/2$, and return the one with minimum total weight. Moreover, to implement this using only comparisons between matchings, we describe a variant of the Bellman-Ford algorithm. For each $u\in L$, let $p_{u}^{(k)}$ be the shortest $M$-alternating path from $L^{\prime}$ to $u$ using at most $k$ matching edges. For matched $u\in L$, let $u^{\prime}\in R$ be its partner. We initialize:
@@
E.3 Generalization to Matroid Intersection
@@
Lemma E.6 (Matching lemmas).
@@
Lemma E.9 (Minimal paths and cycles).
@@
Consider the multiset union of edges $N\cup N^{\prime}\cup 2M$ which takes two copies of every edge from $M$ and a copy of every edge from $N$ and $N^{\prime}$. Since these set of edges enter and leave every vertex exactly twice, the set of edges can be decomposed into a collection of cycles $C_{1},\dots,C_{p}$ for some $p\geq 2$ such that some of the cycles have strictly fewer vertices than $C$. We also have $\sum_{i\in p}\ell(C_{i})=2\ell(C)$. Using Lemma˜E.8(i), we know $\ell(C_{i})\geq 0$ and since $p\geq 2$, there exists a cycle $C_{i}$ with lesser weight or the same weight with lesser number of vertices as $C$ contradicting the minimality of $C$.
@@
    (No re-entry) First, we prove that a minimal path in $\mathcal{P}(y,t)$ never re-enters $X_{1}$ again. If $P=(x_{0},y_{1},x_{1},\dots,y_{k},x_{k},\dots,x_{t-1},y_{t})$ with $x_{k}\in X_{1}$, then observe that there is an edge from $y_{k}$ to $x_{0}$ because $Y+x_{0}\in\mathcal{I}_{1}$ as $x_{0}\in\mathcal{I}_{1}$ which implies $Y+x_{0}-y_{k}\in\mathcal{I}_{1}$ and hence the edge. If we let $C=(x_{0},y_{1},x_{1},\dots,y_{k},x_{0})$, and $P^{\prime}=(x_{k},y_{k+1},\dots,x_{t-1},y_{t})$, we have $\ell(P)=\ell(C)+\ell(P^{\prime})$ implying that $P^{\prime}$ has smaller length and lesser number of arcs than $P$ as $\ell(C)\geq 0$ from Lemma˜E.8(i) contradicting the minimality of $P$. An identical argument can be used to show that vertices are not revisited again in minimal paths.
@@
    Since these set of edges enter and leave every vertex exactly twice, the set of edges can be decomposed into a collection of cycles $C_{1},\dots,C_{p}$ for some $p\geq 2$ such that some of the cycles have strictly fewer vertices than $C$. We also have $\sum\limits_{i\in p}\ell(C_{i})=2\ell(C)$. Using Lemma˜E.8(i), we know $\ell(C_{i})\geq 0$. Consider the two different cycles (say) $C_{1},C_{2}$ that each contain a copy of the edge $(y_{t},x_{0})$. We know that one of these cycles (say) $C_{1}$ should have $\ell(C_{1})\leq\ell(C)$ and $|C_{1}|<|C|$. Removing the $(y_{t},x_{0})$ edge from $C_{1}$ gives a path (say) $P_{1}$ such that $(\ell(P),|P|)\succeq(\ell(P_{1}),|P_{1}|)$ contradicting the minimality of $P$.
@@
Lemma˜E.9 allows us to compare lengths of minimal paths using only comparisons of matroid intersections:
@@
Proof of Theorem˜E.4.
Let $Y_{t}$ be the extreme common independent set of size exactly $t$ starting with $Y_{0}=\emptyset$. Use the modified Bellman-Ford Algorithm to find the shortest augmenting path $P_{t}$ in the exchange graph $H(\mathcal{M}_{1},\mathcal{M}_{2},Y_{t}))$. Using Lemma˜E.8 part (i), the exchange graph has no negative cycles. The shortest path computation takes $O(n^{3})$ comparison queries and time (see section˜E.4). After every update, we know that $Y_{t+1}:=Y_{t}\Delta P_{t}$ is an extreme set using Lemma˜E.8 part (ii). We continue for $n$ steps as long as there is a directed path from $X_{1}$ to $X_{2}$ in the exchange graph. Using Lemma˜E.7, we know that the maximum cardinality intersection is reached when an augmenting path is not present anymore. Finally, we can compare the weights of all the extreme sets obtained to output the minimum weight common intersection $O(n^{4})$ comparisons and $O(n^{4})$ time. ∎
E.4 Shortest paths